### v1

# This is the title

   Abstract: Let $\mathcal{X}$ be a p-adic Hilbert space. Let $A:\mathcal{D}(A)\subseteq\mathcal{X}\to\mathcal{X}$ and $B:\mathcal{D}(B)\subseteq\mathcal{X}\to\mathcal{X}$ be possibly unbounded linear operators. For $x\in\mathcal{D}(A)$ with $\langle x,x\rangle=1$, define $\Delta_{x}(A)\coloneqq\|Ax-\langle Ax,x\rangle x\|.$ Then for all $x\in\mathcal{D}(AB)\cap\mathcal{D}(BA)$ with $\langle x,x\rangle=1$, we show that
$(1)\quad\quad\quad\max\{\Delta_{x}(A),\Delta_{x}(B)\}\geq\frac{\sqrt{\bigg{|}\big{\langle}[A,B]x,x\big{\rangle}^{2}+\big{(}\langle\{A,B\}x,x\rangle-2\langle Ax,x\rangle\langle Bx,x\rangle\big{)}^{2}\bigg{|}}}{\sqrt{|2|}}$
and
$(2)\quad\quad\quad\max\{\Delta_{x}(A),\Delta_{x}(B)\}\geq|\langle(A+B)x,y\rangle|,\quad\forall y\in\mathcal{X}\text{ satisfying }\|y\|\leq 1,\langle x,y\rangle=0.$
We call Inequality (1) as p-adic Heisenberg-Robertson-Schrodinger uncertainty principle and Inequality (2) as p-adic Maccone-Pati uncertainty principle.
Keywords: Uncertainty Principle, p-adic Hilbert space.
Mathematics Subject Classification (2020): 46S10, 12J25, 32P05.
1. Introduction
Let $\mathcal{H}$ be a complex Hilbert space and $A$ be a possibly unbounded self-adjoint linear operator defined on the domain $\mathcal{D}(A)\subseteq\mathcal{H}$. For $h\in\mathcal{D}(A)$ with $\|h\|=1$, define the uncertainty of $A$ at the point $h$ as
$\Delta_{h}(A)\coloneqq\|Ah-\langle Ah,h\rangle h\|=\sqrt{\|Ah\|^{2}-\langle Ah,h\rangle^{2}}.$
In 1929, Robertson [6] derived the following mathematical form of the uncertainty principle of Heisenberg derived in 1927 [3]. Recall that for two operators $A:\mathcal{D}(A)\subseteq\mathcal{H}\to\mathcal{H}$ and $B:\mathcal{D}(B)\subseteq\mathcal{H}\to\mathcal{H}$, we define the commutator $[A,B]\coloneqq AB-BA$ and anti-commutator $\{A,B\}\coloneqq AB+BA$.
Theorem 1.1.
[6, 3, 8, 2] (Heisenberg-Robertson Uncertainty Principle) Let $A:\mathcal{D}(A)\subseteq\mathcal{H}\to\mathcal{H}$ and $B:\mathcal{D}(B)\subseteq\mathcal{H}\to\mathcal{H}$ be self-adjoint linear operators. Then for all $h\in\mathcal{D}(AB)\cap\mathcal{D}(BA)$ with $\|h\|=1$, we have
(1) $\frac{1}{2}\left(\Delta_{h}(A)^{2}+\Delta_{h}(B)^{2}\right)\geq\left(\frac{\Delta_{h}(A)+\Delta_{h}(B)}{2}\right)^{2}\geq\Delta_{h}(A)\Delta_{h}(B)\geq\frac{1}{2}|\langle[A,B]h,h\rangle|.$
In 1930, Schrodinger made the following improvement of Inequality (1).
Theorem 1.2.
[7, 1] (Heisenberg-Robertson-Schrodinger Uncertainty Principle) Let $A:\mathcal{D}(A)\subseteq\mathcal{H}\to\mathcal{H}$ and $B:\mathcal{D}(B)\subseteq\mathcal{H}\to\mathcal{H}$ be self-adjoint linear operators. Then for all $h\in\mathcal{D}(AB)\cap\mathcal{D}(BA)$ with $\|h\|=1$, we have
$\Delta_{h}(A)\Delta_{h}(B)\geq|\langle Ah,Bh\rangle-\langle Ah,h\rangle\langle Bh,h\rangle|$ $=\frac{\sqrt{|\langle[A,B]h,h\rangle|^{2}+|\langle\{A,B\}h,h\rangle-2\langle Ah,h\rangle\langle Bh,h\rangle|^{2}}}{2}$
=($=\frac{\sqrt{(\langle\{A,B\}h,h\rangle-2\langle Ah,h\rangle\langle Bh,h\rangle)^{2}-\langle[A,B]h,h\rangle^{2}}}{2}.$
Surprisingly, in 2014, Maccone and Pati derived the following uncertainty principle which works for any unit vector which is orthogonal to given unit vector [5].
Theorem 1.3.
[5] (Maccone-Pati Uncertainty Principle) Let $A:\mathcal{D}(A)\subseteq\mathcal{H}\to\mathcal{H}$ and $B:\mathcal{D}(B)\subseteq\mathcal{H}\to\mathcal{H}$ be self-adjoint linear operators. Then for all $h\in\mathcal{D}(A)\cap\mathcal{D}(B)$ with $\|h\|=1$, we have
$\Delta_{h}(A)^{2}+\Delta_{h}(B)^{2}\geq\frac{1}{2}\left(|\langle(A+B)h,k\rangle|^{2}+|\langle(A-B)h,k\rangle|^{2}\right),\quad\forall k\in\mathcal{H}\text{ satisfying }\|k\|=1,\langle h,k\rangle=0.$
As the study of p-adic Hilbert spaces is equally important as the study of Hilbert spaces, we naturally ask the following question.
Question 1.4.
What are p-adic versions of Theorems 1.2 and 1.3?
In this paper, we answer Question 1.4.
2. p-adic Heisenberg-Robertson-Schrodinger Uncertainty Principle and p-adic Maccone-Pati uncertainty principle
We are going to consider the following notion of p-adic Hilbert space which is introduced by Kalisch [4] in 1947.
Definition 2.1.
[…]
  1. (i)
    If $x\in\mathcal{X}$ is such that $\langle x,y\rangle=0$ for all $y\in\mathcal{X}$, then $x=0$.
  2. (ii)
    $\langle x,y\rangle=\langle y,x\rangle$ for all $x,y\in\mathcal{X}$.
  3. (iii)
    $\langle x,\alpha y+z\rangle=\alpha\langle x,y\rangle+\langle x,z\rangle$ for all $\alpha\in\mathbb{K}$, for all $x,y,z\in\mathcal{X}$.
  4. (iv)
    $|\langle x,y\rangle|\leq\|x\|\|y\|$ for all $x,y\in\mathcal{X}$.
Following are standard examples.
Example 2.2.
Let $d\in\mathbb{N}$ and $\mathbb{K}$ be a non-Archimedean valued field. Then $\mathbb{K}^{d}$ is a p-adic Hilbert space w.r.t. norm
$\|(x_{j})_{j=1}^{d}\|\coloneqq\max_{1\leq j\leq d}|x_{j}|,\quad\forall(x_{j})_{j=1}^{d}\in\mathbb{K}^{d}$
and p-adic inner product
$\langle(x_{j})_{j=1}^{d},(y_{j})_{j=1}^{d}\rangle\coloneqq\sum_{j=1}^{d}x_{j}y_{j},\quad\forall(x_{j})_{j=1}^{d},(y_{j})_{j=1}^{d}\in\mathbb{K}^{d}.$
Example 2.3.
Let $\mathbb{K}$ be a non-Archimedean valued field. Define
$c_{0}(\mathbb{N},\mathbb{K})\coloneqq\{(x_{n})_{n=1}^{\infty}:x_{n}\in\mathbb{K},\forall n\in\mathbb{N},\lim_{n\to\infty}x_{n}=0\}.$
Then $c_{0}(\mathbb{N},\mathbb{K})$ is a p-adic Hilbert space w.r.t. norm
$\|(x_{n})_{n=1}^{\infty}\|\coloneqq\sup_{n\in\mathbb{N}}|x_{n}|,\quad\forall(x_{n})_{n=1}^{\infty}\in c_{0}(\mathbb{N},\mathbb{K})$
and p-adic inner product
$\langle(x_{n})_{n=1}^{\infty},(y_{n})_{n=1}^{\infty}\rangle\coloneqq\sum_{n=1}^{\infty}x_{n}y_{n},\quad\forall(x_{n})_{n=1}^{\infty},(y_{n})_{n=1}^{\infty}\in c_{0}(\mathbb{N},\mathbb{K}).$
Let $\mathcal{X},\mathcal{Y}$ be p-adic Hilbert spaces and $T:\mathcal{X}\to\mathcal{Y}$ be a linear operator. We say that $T$ is adjointable if there is a linear operator, denoted by $T^{*}:\mathcal{Y}\to\mathcal{X}$ such that $\langle Tx,y\rangle=\langle x,T^{*}y\rangle$, $\forall x\in\mathcal{X},\forall y\in\mathcal{X}$. Note that (i) in Definition 2.1 says that adjoint, if exists, is unique. An adjointable linear operator $T:\mathcal{X}\to\mathcal{X}$ is said to be self-adjoint if $T^{*}=T$.
Let $A$ be a possibly unbounded linear operator (need not be self-adjoint) defined on domain $\mathcal{D}(A)\subseteq\mathcal{X}$. For $x\in\mathcal{D}(A)$ with $\langle x,x\rangle=1$, define the uncertainty of $A$ at the point $x$ as
$\Delta_{x}(A)\coloneqq\|Ax-\langle Ax,x\rangle x\|.$
We now have the p-adic version of Theorem 1.2.
Theorem 2.4.
(p-adic Heisenberg-Robertson-Schrodinger Uncertainty Principles) Let $\mathcal{X}$ be a p-adic Hilbert space. Let $A:\mathcal{D}(A)\subseteq\mathcal{X}\to\mathcal{X}$ and $B:\mathcal{D}(B)\subseteq\mathcal{X}\to\mathcal{X}$ be linear operators. Then for all $x\in\mathcal{D}(AB)\cap\mathcal{D}(BA)$ with $\langle x,x\rangle=1$, we have
  1. (i)
    $\Delta_{x}(A)\Delta_{x}(B)\geq|\langle Ax,Bx\rangle-\langle Ax,x\rangle\langle Bx,x\rangle|=|\langle Bx,Ax\rangle-\langle Ax,x\rangle\langle Bx,x\rangle|.$
    In particular, if $A$ and $B$ are self-adjoint, then
    $\Delta_{x}(A)\Delta_{x}(B)\geq|\langle BAx,x\rangle-\langle Ax,x\rangle\langle Bx,x\rangle|=|\langle ABx,x\rangle-\langle Ax,x\rangle\langle Bx,x\rangle|.$
  2. (ii)
    $\max\{\Delta_{x}(A),\Delta_{x}(B)\}\geq\frac{\sqrt{\bigg{|}\big{\langle}[A,B]x,x\big{\rangle}^{2}+\big{(}\langle\{A,B\}x,x\rangle-2\langle Ax,x\rangle\langle Bx,x\rangle\big{)}^{2}\bigg{|}}}{\sqrt{|2|}}.$
  3. (iii)
    If $A$ and $B$ are adjointable, then
    $\max\{\Delta_{x}(A),\Delta_{x}(B)\}\geq\sqrt{\left|\langle(A^{*}A+B^{*}B)x,x\rangle-\frac{\langle(A+B)x,x\rangle^{2}+\langle(A-B)x,x\rangle^{2}}{2}\right|}.$
    In particular, if $A$ and $B$ are self-adjoint, then
    $\max\{\Delta_{x}(A),\Delta_{x}(B)\}$ $\geq\sqrt{\left|\langle(A^{2}+B^{2})x,x\rangle-\frac{\langle(A+B)x,x\rangle^{2}+\langle(A-B)x,x\rangle^{2}}{2}\right|}$
    $=\sqrt{\left|\frac{\langle(A+B)^{2}x,x\rangle+\langle(A-B)^{2}x,x\rangle-\langle(A+B)x,x\rangle^{2}-\langle(A-B)x,x\rangle^{2}}{2}\right|}.$
  4. (iv)
    If $A$ and $B$ are adjointable, then
    $\max\{\Delta_{x}(A),\Delta_{x}(B)\}\geq\sqrt{|\langle(A^{*}A-B^{*}B)x,x\rangle-\langle(A+B)x,x\rangle\langle(A-B)x,x\rangle|}.$
    In particular, if $A$ and $B$ are self-adjoint, then
    $\max\{\Delta_{x}(A),\Delta_{x}(B)\}\geq\sqrt{|\langle(A^{2}-B^{2})x,x\rangle-\langle(A+B)x,x\rangle\langle(A-B)x,x\rangle|}.$
  5. (v)
    $\max\{\Delta_{x}(A),\Delta_{x}(B)\}\geq\sqrt{|\langle(A+B)x,(A+B)x\rangle-\langle(A+B)x,x\rangle^{2}|}.$
  6. (vi)
    $\max\{\Delta_{x}(A),\Delta_{x}(B)\}\geq\sqrt{|\langle(A-B)x,(A-B)x\rangle-\langle(A-B)x,x\rangle^{2}|}.$
Proof.
Let $x\in\mathcal{D}(AB)\cap\mathcal{D}(BA)$ be such that $\langle x,x\rangle=1$.
  1. (i)
    By using the definition of p-adic inner product,
    $\Delta_{x}(A)\Delta_{x}(B)$ $=\|Ax-\langle Ax,x\rangle x\|\|Bx-\langle Bx,x\rangle x\|$
    $\geq|\langle Ax-\langle Ax,x\rangle x,Bx-\langle Bx,x\rangle x\rangle|$
    $=|\langle Ax,Bx\rangle-\langle Ax,x\rangle\langle Bx,x\rangle|.$
  2. (ii)
    By making a direct expansion and simplification, we see that
    $\big{\langle}[A,B]x,x\big{\rangle}^{2}+\big{(}\langle\{A,B\}x,x\rangle-2\langle Ax,x\rangle\langle Bx,x\rangle\big{)}^{2}$
    $=\big{(}\langle ABx,x\rangle-\langle BAx,x\rangle\big{)}^{2}+\big{\langle}\{A,B\}x,x\big{\rangle}^{2}+4\langle Ax,x\rangle^{2}\langle Bx,x\rangle^{2}-4\langle\{A,B\}x,x\rangle\langle Ax,x\rangle\langle Bx,x\rangle$
    $=\big{(}\langle ABx,x\rangle-\langle BAx,x\rangle\big{)}^{2}+\big{(}\langle ABx,x\rangle+\langle BAx,x\rangle\big{)}^{2}+4\langle Ax,x\rangle^{2}\langle Bx,x\rangle^{2}$
    $\quad-4\langle ABx,x\rangle\langle Ax,x\rangle\langle Bx,x\rangle-4\langle BAx,x\rangle\langle Ax,x\rangle\langle Bx,x\rangle$
    $=2\langle ABx,x\rangle^{2}+2\langle BAx,x\rangle^{2}+4\langle Ax,x\rangle^{2}\langle Bx,x\rangle^{2}-4\langle ABx,x\rangle\langle Ax,x\rangle\langle Bx,x\rangle-4\langle BAx,x\rangle\langle Ax,x\rangle\langle Bx,x\rangle$
    $=2(\langle ABx,x\rangle-\langle Ax,x\rangle\langle Bx,x\rangle)^{2}+2(\langle BAx,x\rangle-\langle Ax,x\rangle\langle Bx,x\rangle)^{2}.$
    Therefore
    $\left|\big{\langle}[A,B]x,x\big{\rangle}^{2}+\big{(}\langle\{A,B\}x,x\rangle-2\langle Ax,x\rangle\langle Bx,x\rangle\big{)}^{2}\right|$
    $=|2|\left|(\langle ABx,x\rangle-\langle Ax,x\rangle\langle Bx,x\rangle)^{2}+(\langle BAx,x\rangle-\langle Ax,x\rangle\langle Bx,x\rangle)^{2}\right|$
    $\leq|2|\max\left\{\left|(\langle ABx,x\rangle-\langle Ax,x\rangle\langle Bx,x\rangle)^{2}\right|,\left|(\langle BAx,x\rangle-\langle Ax,x\rangle\langle Bx,x\rangle)^{2}\right|\right\}$
    $=|2|\max\left\{\left|\langle ABx,x\rangle-\langle Ax,x\rangle\langle Bx,x\rangle\right|^{2},\left|\langle BAx,x\rangle-\langle Ax,x\rangle\langle Bx,x\rangle\right|^{2}\right\}$
    $\leq|2|\max\left\{\max\{\Delta_{x}(A)^{2},\Delta_{x}(B)^{2}\},\max\{\Delta_{x}(B)^{2},\Delta_{x}(A)^{2}\}\right\}=|2|\max\{\Delta_{x}(A)^{2},\Delta_{x}(B)^{2}\}.$
  3. (iii)
    By using the non-Archimedean triangle inequality and the definition of p-adic inner product,
    $\max\{\Delta_{x}(A),\Delta_{x}(B)\}$ $=\max\{\|Ax-\langle Ax,x\rangle x\|,\|Bx-\langle Bx,x\rangle x\|\}$
    $\geq\max\{\sqrt{|\langle Ax-\langle Ax,x\rangle x,Ax-\langle Ax,x\rangle x\rangle|},\sqrt{|\langle Bx-\langle Bx,x\rangle x,Bx-\langle Bx,x\rangle x\rangle|}\}$
    $=\max\{\sqrt{|\langle Ax,Ax\rangle-\langle Ax,x\rangle^{2}|},\sqrt{|\langle Bx,Bx\rangle-\langle Bx,x\rangle^{2}|}\}$
    $=\sqrt{\max\{|\langle Ax,Ax\rangle-\langle Ax,x\rangle^{2}|,|\langle Bx,Bx\rangle-\langle Bx,x\rangle^{2}|\}}$
    $\geq\sqrt{|\langle Ax,Ax\rangle-\langle Ax,x\rangle^{2}+\langle Bx,Bx\rangle-\langle Bx,x\rangle^{2}|}$
    $=\sqrt{|\langle A^{*}Ax,x\rangle+\langle B^{*}Bx,x\rangle-(\langle Ax,x\rangle^{2}+\langle Bx,x\rangle^{2})|}$
    $=\sqrt{\left|\langle(A^{*}A+B^{*}B)x,x\rangle-\frac{\langle(A+B)x,x\rangle^{2}+\langle(A-B)x,x\rangle^{2}}{2}\right|}.$
  4. (iv)
    Using initial calculations in (iii),
    $\max\{\Delta_{x}(A),\Delta_{x}(B)\}$ $\geq\sqrt{\max\{|\langle Ax,Ax\rangle-\langle Ax,x\rangle^{2}|,|\langle Bx,Bx\rangle-\langle Bx,x\rangle^{2}|\}}$
    $\geq\sqrt{|\langle Ax,Ax\rangle-\langle Ax,x\rangle^{2}-\langle Bx,Bx\rangle+\langle Bx,x\rangle^{2}|}$
    $=\sqrt{|\langle A^{*}Ax,x\rangle-\langle B^{*}Bx,x\rangle-(\langle Ax,x\rangle^{2}-\langle Bx,x\rangle^{2})|}$
    $=\sqrt{|\langle(A^{*}A-B^{*}B)x,x\rangle-\langle(A+B)x,x\rangle\langle(A-B)x,x\rangle|}.$
  5. (v)
    Using ultrametric inequality first and then using p-adic inner product we get
    $\max\{\Delta_{x}(A),\Delta_{x}(B)\}$ $=\max\{\|Ax-\langle Ax,x\rangle x\|,\|Bx-\langle Bx,x\rangle x\|\}$
    $\geq\|Ax-\langle Ax,x\rangle x+Bx-\langle Bx,x\rangle x\|$
    $\geq\sqrt{|\langle Ax-\langle Ax,x\rangle x+Bx-\langle Bx,x\rangle x,Ax-\langle Ax,x\rangle x+Bx-\langle Bx,x\rangle x\rangle|}$
    $=\sqrt{|\langle Ax,Ax\rangle+\langle Bx,Bx\rangle+2\langle Ax,Bx\rangle-2\langle Ax,x\rangle\langle Bx,x\rangle-\langle Ax,x\rangle^{2}-\langle Bx,x\rangle^{2}|}$
    $=\sqrt{|\langle(A+B)x,(A+B)x\rangle-\langle(A+B)x,x\rangle^{2}|}.$
  6. (vi)
    Using initial calculations in (v),
    $\max\{\Delta_{x}(A),\Delta_{x}(B)\}$ $=\max\{\|Ax-\langle Ax,x\rangle x\|,\|Bx-\langle Bx,x\rangle x\|\}$
    $\geq\|Ax-\langle Ax,x\rangle x-Bx+\langle Bx,x\rangle x\|$
    $\geq\sqrt{|\langle Ax-\langle Ax,x\rangle x-Bx+\langle Bx,x\rangle x,Ax-\langle Ax,x\rangle x-Bx+\langle Bx,x\rangle x\rangle|}$
    $=\sqrt{|\langle(A-B)x,(A-B)x\rangle-\langle(A-B)x,x\rangle^{2}|}.$
Note that for self-adjoint operators $A$ and $B$, we have
$\langle[A,B]x,x\rangle=\langle ABx,x\rangle-\langle BAx,x\rangle=\langle Bx,Ax\rangle-\langle Ax,Bx\rangle=\langle Bx,Ax\rangle-\langle Bx,Ax\rangle=0$
and
$\langle\{A,B\}x,x\rangle$ $=\langle ABx,x\rangle+\langle BAx,x\rangle=\langle Bx,Ax\rangle+\langle Ax,Bx\rangle=\langle Bx,Ax\rangle+\langle Bx,Ax\rangle$
$=2\langle Bx,Ax\rangle=2\langle ABx,x\rangle.$
We next derive p-adic version of Theorem 1.3.
Theorem 2.5.
(p-adic Maccone-Pati Uncertainty Principle) Let $\mathcal{X}$ be a p-adic Hilbert space. Let $A:\mathcal{D}(A)\subseteq\mathcal{X}\to\mathcal{X}$ and $B:\mathcal{D}(B)\subseteq\mathcal{X}\to\mathcal{X}$ be linear operators. Then for all $x\in\mathcal{D}(A)\cap\mathcal{D}(B)$ with $\langle x,x\rangle=1$, we have
$\max\{\Delta_{x}(A),\Delta_{x}(B)\}\geq|\langle(A+B)x,y\rangle|,\quad\forall y\in\mathcal{X}\text{ satisfying }\|y\|\leq 1,\langle x,y\rangle=0$
and
$\max\{\Delta_{x}(A),\Delta_{x}(B)\}\geq|\langle(A-B)x,y\rangle|,\quad\forall y\in\mathcal{X}\text{ satisfying }\|y\|\leq 1,\langle x,y\rangle=0.$
Proof.
Let $x\in\mathcal{D}(A)\cap\mathcal{D}(B)$ be such that $\langle x,x\rangle=1$. Let $y\in\mathcal{X}$ satisfies $\|y\|\leq 1$ and $\langle x,y\rangle=0.$ Then
$|\langle(A+B)x,y\rangle|$ $=|\langle Ax-\langle Ax,x\rangle x+Bx-\langle Bx,x\rangle x,y\rangle|$
$\leq\|Ax-\langle Ax,x\rangle x+Bx-\langle Bx,x\rangle x\|\|y\|$
$\leq\|Ax-\langle Ax,x\rangle x+Bx-\langle Bx,x\rangle x\|$
$\leq\max\{\|Ax-\langle Ax,x\rangle x\|,\|Bx-\langle Bx,x\rangle x\|\}$
$=\max\{\Delta_{x}(A),\Delta_{x}(B)\}.$
3. Acknowledgments
This paper has been partially developed when the author attended the workshop “Quantum groups, tensor categories and quantum field theory”, held in University of Oslo, Norway from January 13 to 17, 2025. This event was organized by University of Oslo, Norway and funded by the Norwegian Research Council through the “Quantum Symmetry” project.